You will be given a series of questions regarding a normal distribution, you will be asked to either determine the percentage above or below particular raw scores; or to calculate the raw score that will correspond to a particular percentage.

You will be asked to calculate either raw scores or percentages. For each question write out your calculation, the appropriate Z score, what on the curve you should be shading, the exact percentage from the normal curve table and the final answer.
A group of students take a Statistics Exam where the average was M = 90 and the standard deviation was SD = 7.8. Answer the following questions regarding this distribution using your normal curve table. Depending on the problem, be sure to identify the raw scores, Z scores, what in particular you shaded when creating a curve and the final answer. (5 pts each)
1.​If you score a 99, what percentage of the population scored above you?
2.​What raw score is needed to be in the top 8%?
3.​What raw score is needed to reach the 14th percentile?
4.​What range of raw scores makes up the middle 40%?
5.​If you score an 80, what percentage of the population scored above you?
6.​What percentage of the population lands between the raw scores of 82 and 107?
7.​If you score a 85, what percentage of the population scored above you?
8.​What raw score is needed to be in the bottom 16%?
9.​If you score a 96, what percentage of the population scored above you?
10.​What raw score is needed to be in the top 17%?
11.​What raw score is needed to reach the 90th percentile?
12.​What percentage of the population lands between the raw scores of 83 and 100?
13.​If you score a 85, what percentage of the population below you?
14.​What raw score is needed to be in the bottom 28%?
15.​What percentage of the population lands between the raw scores of 86 and 105?

TABLE B:1 THE DISTRIBUTION ol A ū 2013 SE WD WO 1990 wa B1
Answer #1

normal ns: Marks in a statistics exam follow distribution with mean (1M) = 90 S.D(E)= 7.8 tet x=99, Then Pro (X >99) = ? Fo6=2) Top 8%: 92% Let x be the raw score that is the least in top 8%. To the left of this y, there is 100-8=927 of population. NFor 82, Z = X-4 _ 82-90 = -1.03. 7:8 for 107, z = 107-90 = 2.18 7.8 from the above came, it is clear that Pro (82 < x <107) =